A radioactive sample of thallium-207 has a half-life of 4.77 minutes has an activity of 1500 Ci.
What is the mass of the sample of thallium? (to 2 s.f and in µg)
1
Expert's answer
2020-06-01T14:18:24-0400

Activity of the sample is "R=1500 Ci=1500(3.7*10^{10})"

"=5.55*10^{13} decays\/s"

The expression for the activity of the sample is

"R= \frac{ln2 N_A m}{M T_{1/2}}=\frac{0.693 N_A m}{M T_{1/2}}"

Here "N_A" Avogadro number.

"m" is the mass of sample.

"M" is the molecular mass.

Isolate the above equation for "m" ,

"m=\frac{RM T_{1/2}}{0.693 N_A}"

"=\frac{(5.55*10^{13})(206.977*10^{-3})(4.77*60)}{0.693 (6.022*10^{23})}"

"=7.99*10^{-9} kg=7.99*10^{-6} g =7.99 \mu g"